OpenStudy (anonymous):

In an egg carton there are 12 eggs, of which 9 are hard-boiled and 3 are raw. Six of the eggs are chosen at random to take to a picnic (yes, the draws are made without replacement). Find the chance that at least one of the chosen eggs is raw.

OpenStudy (shubhamsrg):

find probab. that none of the eggs are raw and subtract it from 1 .

OpenStudy (shubhamsrg):

any attempt so far ?

OpenStudy (perl):

X is a hypergeometric random variable , P ( X > = 1 ) = 1 - P ( X < 1) = 1 - P ( X=0)

OpenStudy (perl):